GROUPS ST ANDREWS 1997 IN BATH

This conference will be held at the University of Bath, Bath, England from Saturday 26 July 1997 (arrival day) to Saturday 9 August 1997 (departure day). It will be a large international conference and, based on experience from previous Groups St Andrews conferences, we expect around 350 participants.

The main speakers, each giving a course of about four lectures, will be

• Laszlo Babai, University of Chicago

• Martin Bridson, Pembroke College, Oxford

• Chris Brookes, Corpus Christi College, Cambridge

• Cheryl Praeger, University of Western Australia

• Aner Shalev, The Hebrew University of Jerusalem

The courses by the main speakers will be in the first week of the conference. The second week will include a programme of invited one hour talks. There will be a full seminar programme in both weeks of the conference.

Bath is an attractive and historic town easily reached from London Heathrow airport. As with previous Groups St Andrews conferences there will be an interesting social programme. Accompanying relatives, families and friends will be most welcome.

Further information about the conference and the city of Bath may be viewed at the WWW page: http://www.bath.ac.uk/∼masgcs/gps97/

Conference email address: groups97@dcs.st-and.ac.uk

If you wish to receive further information about the conference, when available, including an application form, please reply by email to groups97@dcs.stand.ac.uk or complete the form below and send it to

Groups St Andrews 1997 in Bath Mathematical Institute North Haugh St Andrews Fife KY16 9SS Scotland

Colin Campbell, Edmund Robertson, Geoff Smith (organizing committee)

I am interested in receiving further information about

Groups St Andrews 1997 in Bath

taking place from 26 July 1997 to 9 August 1997. Title: Name: Preferred first name: Postal Address:

email address: Comments: